OpenStudy (anonymous):

prove this is right with identities: (sin a)^4 + (cos a)^4 = 1 - 2 * (sin a)^2 * (cos a)^2

OpenStudy (anonymous):

\[\sin ^{4}\alpha + \cos ^{4}\alpha = 1-2 \sin ^{2}\alpha \cos ^{2}\alpha\] please help

hartnn (hartnn):

add and subtract 2 * (sin a)^2 * (cos a)^2 on left side : (sin a)^4 + (cos a)^4

hartnn (hartnn):

=[(sin a)^4 + (cos a)^4+2 * (sin a)^2 * (cos a)^2]-2 * (sin a)^2 * (cos a)^2 did you get what i did ?

OpenStudy (anonymous):

trying to... :) thank you...

hartnn (hartnn):

then notice that\( (\sin a)^4 =(\sin^2a)^2 \) same for cos^4 a and that [(sin a)^4 + (cos a)^4+2 * (sin a)^2 * (cos a)^2] is of the form \(x^2+y^2+2xy=(x+y)^2\)

hartnn (hartnn):

are u getting any of this ? please interact and ask doubts....

OpenStudy (anonymous):

i'm writing it down and trying to get to eah of your steps by myself so it takes some times...thank u man :)

hartnn (hartnn):

good :) also think of next steps...

OpenStudy (anonymous):

okay, got it... sin^2a + cos^2a = 1 1^2=1 you're awesome :) thanks alot

hartnn (hartnn):

welcome ^_^ thanks for the compliment :)